OpenStudy (anonymous):

I have a few questions I need to be checked and one I need to go over, anyone up for the task? i'll post the questions down below...

OpenStudy (anonymous):

Maria brings d dollars to an amusement park. The park charges an admission fee of $20. Each game at the park costs $1.50. Which inequality shows the number of games, g, which Maria can play? (4 points) 20 + 1.5g ≥ d 20 + 1.5g < d 20 + 1.5g > d 20 + 1.5g ≤ d<<<My Answer 7. Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2 8. A company makes jewelry. It costs $6,400 to make 10 pieces of jewelry and $11,200 to make 20 pieces of jewelry. Which equation models the cost, C(x), as a linear function of the number of pieces of jewelry, x? C(x) = 480x + 1600<<< My Answer C(x) = 1600x + 480 C(x) = 480x − 1600 C(x) = 1600x − 480

OpenStudy (amistre64):

if you can do 8, you can do 7

OpenStudy (anonymous):

I had help with 8, but I still don't understand how to solve it, he tried to explain but got frustrated and just gave me the answer and left... do you think you could help explain it better

OpenStudy (amistre64):

you have 4 equations to "test" with; and values for x and y to plug in; see what fits

OpenStudy (amistre64):

or, go the long route of coming up with an actual equation from the given data

OpenStudy (anonymous):

okay, im the process of doing so... I just need a little to check the all... why couldnt the guy just say that before... lol he made it seem much more complicated

OpenStudy (amistre64):

if a test gives you options, the quickest route is to see what fits; if you have to come up with an equation from scratch, the longer way is required; but this here just amounts to plugging in the values and see what fits

OpenStudy (anonymous):

so, the answer would be B, right?

OpenStudy (amistre64):

lets see :) 2x + y = 2 2 -2 ----------- 4 - 2 = 2 , that ones good 2x + y = 2 -5 12 ----------- -10+12=2 , i like it

OpenStudy (anonymous):

yay! :) Thanks! I do have two more that are confusing me, could you help with them as well?

OpenStudy (amistre64):

post them to the left so that others can participate and get just rewards as well.

OpenStudy (anonymous):

oops I didnt mean to do that

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I deleted it, no worries, ill go post the other two now, thanks again you guys!! :)